# 1]Centre Étude et simulation du climat à l’échelle régionale (ESCER), Département des Sciences de la Terre et de l’Atmosphère, Université du Québec à Montréal, Montréal, Québec, Canada

### v2

@@ Abstract
The 2021 Pacific Northwest (PN2021) heatwave shattered historical temperature records, culminating in Lytton, Canada’s unprecedented 49.6C observation—a 4.6C increase over the country’s previous record measurement (White et al., 2023; Mass et al., 2024). This event, virtually implausible under preindustrial conditions (Philip et al., 2022), exemplifies a critical challenge in climate science: determining the upper bounds of what is physically possible for different weather extremes under current or future climatic conditions.
The PN2021 heatwave emerged from persistent atmospheric blocking sustained by large-scale Rossby waves that disrupted zonal flow and stalled a high-pressure system over the region (Mass et al., 2024). These low-wavenumber waves suppressed cloud formation and prolonged subsidence, adiabatically warming near-surface air masses. Concurrent dry soil conditions, amplified by antecedent spring drought, reduced latent heat flux and increased sensible heat, further intensifying surface temperatures (Mass et al., 2024). The interplay of the blocking persistence and land-atmosphere feedbacks culminated in the unprecedented temperature anomalies.
@@
To systematically explore such extremes, storylines are increasingly used, representing physically consistent sequences of weather events that depict how a counterfactual extreme event might occur (Hazeleger et al., 2015; Shepherd, 2019; Sillmann et al., 2021). This approach enables a mechanistic exploration of how minor perturbations can lead to the amplification of extreme events. Here, we use a novel differentiable modeling framework to demonstrate that targeted initial-condition perturbations can further amplify these typical extreme trajectories, giving extreme heatwave storylines.
@@
Identifying storylines for the most extreme weather events is a needle-in-a-haystack problem due to their inherent rarity. The traditional approach is the use of single-model initial-condition large ensembles (Deser et al., 2020; Suarez-Gutierrez et al., 2020; Maher et al., 2021; Diffenbaugh and Davenport, 2021), and more recently, so-called huge ensembles (Mahesh et al., 2024a, b), in which vast numbers of model runs allow the exploration of a wide range of potential outcomes. By systematically increasing ensemble size, the chances of capturing low-probability extremes increase. However, these ensembles are computationally demanding and not very effective at sampling the full range of outcomes. In addition, due to their high computational cost, it is virtually impossible to perform such ensembles using kilometer-scale simulations, which are required to well simulate some types of weather extreme events (e.g., extreme convective precipitation).
@@
In recent years, a number of approaches have been proposed to generate extreme event storylines (Ragone et al., 2018; Plotkin et al., 2019; Webber et al., 2019; Yiou and Jézéquel, 2020; Gessner et al., 2021; Fischer et al., 2023). These methods focus computing resources on specific extreme events, instead of continuous long simulations. Some approaches enhance the likelihood of simulating extreme events by constructing targeted ensembles (Ragone et al., 2018; Webber et al., 2019; Fischer et al., 2023). Fischer et al. (2023) focus on generating an initial condition ensemble of climate model simulations of known extreme events using a method named ensemble boosting. They applied this approach to the PN2021 heatwave and by perturbing the initial conditions using numerical noise for 500 members, they found a 5-day running average of daily maximum temperature anomalies up to 2.9C larger than the unperturbed event. Other approaches to construct storylines of extreme events use the large deviation algorithm (Ragone et al., 2018; Ragone and Bouchet, 2021; Noyelle et al., 2025) where an ensemble of simulations is ran and members are periodically pruned or cloned such that an ensemble most likely to lead to an extreme event is generated. Focusing on western European heatwaves and using the large deviation algorithm, Ragone et al. (2018) generated an ensemble which has a mean 2C anomaly compared to a control ensemble of 128 members. Other applications of the algorithm showed its ability to identify even more extreme events, with ensembles with mean anomalies of 4C (Ragone and Bouchet, 2021). Meanwhile, Plotkin et al. (2019) introduced a variational data assimilation technique, optimized with a 4D-Var inspired method to intensify past extreme tropical cyclones with minimal perturbations. This approach is closely related to the method presented here; however, we leverage automatic differentiation and computationally efficient ML-based models.
@@
Recent advances in machine learning (ML) have led to the development of transformative tools for weather and climate modeling. Neural network architectures like GraphCast (Lam et al., 2023), Pangu-Weather (Bi et al., 2023), FourCastNet (Pathak et al., 2022; Kurth et al., 2023), and FuXi (Chen et al., 2023) have demonstrated forecasting skill comparable to that of traditional numerical weather prediction systems, but at significantly reduced computational costs (Rasp et al., 2024; Pasche et al., 2025; Ennis et al., 2025). In addition to their reduced computational costs, these models by construction allow us to define optimization problems on them that can be solved through gradient-based optimizers. Rasp et al. (2024) introduced a standardized benchmark to compare the various ML models against ERA5 and the European Centre for Medium-Range Weather Forecasts’s (ECMWF) integrated forecast system (IFS). Using this benchmark, it is shown that deterministic, data-driven methods such as Pangu-Weather, GraphCast, and FuXi result in similar root-mean-square error (RMSE) in forecasting near-surface temperature, wind, and pressure up to ten days ahead. However, their forecast skill deteriorates rapidly for longer lead times, resulting in overly smoothed predictions. Using three case studies, Pasche et al. (2025) evaluated GraphCast, Pangu-Weather and FourCastNet against ERA5 reanalysis and ECMWF’s IFS for the PN2021 heatwave, the 2023 South Asian humid heatwave and a 2021 North American winter storm. They find that all data-driven models systematically underestimate the peak 2m temperature during the PN2021 heatwave, with root mean square error (RMSE) values at grid points near Vancouver, Seattle, and Portland exceeding twice the 10-day IFS error and reaching up to four times that value in Portland. During the South Asian humid heatwave, data-driven forecasts of heat index computed from 2-m air temperature and 1000-hPa relative humidity underpredicted observed peaks more strongly than IFS, particularly over Bangladesh. For the North American winter storm, data-driven forecasts of wind chill at College Station, Texas, achieved lower peak errors than IFS, with Pangu-Weather and GraphCast outperforming the operational model.
@@
An alternative to purely data-driven approaches is the use of hybrid models, such as NeuralGCM (Kochkov et al., 2024), which combines a traditional dynamical core with ML components. (Duan et al., 2025) have shown NeuralGCM’s ability to hindcast the PN2021 heatwave, though due to the lack of processes (such as land-atmosphere feedbacks), the intensity of the heatwaves tends to be underestimated. Similarly to the purely data-driven models, NeuralGCM produces surface variables forecasts with skill comparable to that of the ECMWF IFS system Rasp et al. (2024). Moreover, the use of the dynamical core both prevents the evolved fields from being overly smoothed and enhances numerical stability. These benefits allow for longer time integrations and make the model suitable for climate studies (Kochkov et al., 2024).
@@
These new types of models are by construction differentiable through automatic differentiation Gelbrecht et al. (2023). The automatic differentiation property enables efficient optimization, allowing gradient-based exploration storylines in high-dimensional climate models. This is in line with many new extreme event opportunities enabled by ML models (Materia et al., 2024; Camps-Valls et al., 2025). Leveraging automatic differentiation, recent studies have implemented variational data assimilation techniques using neural networks, with applications ranging from toy models, such as the Lorenz 96 system (Lorenz, 1995), to reduced-order physical representations of the atmosphere (Solvik et al., 2025; Manshausen et al., 2024). Additionally, Vonich and Hakim (2024) demonstrated that the differentiability of GraphCast allows for a more accurate reconstruction of the initial conditions that led to the PN2021 heatwave compared to using ERA5 reanalysis data. Baño-Medina et al. (2025) explores the use of ML models and automatic differentiation to perform sensitivity analysis of the initial conditions leading to the development of cyclone Xynthia. Their findings suggest that gradients computed from the data-driven weather model at a 36-hour lead time exhibit sensitivity structures that closely resemble those generated by the adjoint of a dynamical model. In other words, the evolved perturbations from both approaches lead to similar impacts on the cyclone’s evolution.
@@
In this study, we focus on the PN2021 heatwave event due to its well-documented synoptic drivers, including persistent blocking patterns and quasi-resonant wave amplification (Mass et al., 2024), and its prevalence in extreme event studies (Lucarini et al., 2023; Fischer et al., 2023; White et al., 2023; Philip et al., 2022). We use the automatic differentiation feature of the NeuralGCM model to optimize perturbed initial conditions, and we identify trajectories where enhanced geopotential height anomalies intensify downstream near-surface temperature extremes. These storylines reveal heatwave intensity increases of $3.7^{\circ}$ beyond the extreme temperatures obtained from a 75-member ensemble run using NeuralGCM for the event, analogous to the ensemble boosting approach (Fischer et al., 2023). Our results demonstrate the potential of differentiable hybrid models for investigating worst-case scenarios, offering a computationally efficient alternative to traditional, computationally expensive, large ensembles.
@@ 1.1 Initial Conditions Optimization Problem
with $b$ representing some nonlinear operator, $t$ the time and $\vec{x}_{0}$ some initial state. The solution is given by $\vec{X}(t)=S_{t}\vec{x}_{0}$ where $S_{t}$ is the evolution operator up to time $t$. The core aim of the optimization problem is to identify the initial conditions $\vec{x}_{0}$ that drive the dynamical system toward an extreme desired state, represented by a target observable $\mathcal{O}(\vec{X}(t))$. Given a baseline initial state $\vec{x}_{0}^{b}$, we define a perturbation $\Delta\vec{x}_{0}^{i}=\vec{x}_{0}^{b}-\vec{x}_{0}^{i}$. The optimization problem is formulated as minimizing a loss function:
@@
In particular, we pick our observable to be the temperature over a domain $\mathcal{D}$ and over a period of time $\tau$ at the 1000-hPa pressure level of the model. Multiple functions can be considered to achieve this, but our main results use :
[…]
@@
where $\mathcal{D}$ corresponds to the region shown in Fig. 4, and $\tau$ is set to 5 days. The terms in the objective function are normalized by their initial means, with $T_{\text{ref}}$ representing a characteristic temperature scale and $\Delta x_{i,\text{ref}}$ denoting a reference perturbation scale for each component $i$. The normalization scale for each perturbed variable, $\Delta x_{ref,i}$, is defined as the absolute mean of each respective initial field. This ensures each terms is of a similar magnitudes.
1.2 Numerical Implementation using NeuralGCM
To simulate the dynamics and evaluate the loss function, we use the NeuralGCM model (Kochkov et al., 2024) with a horizontal grid spacings of $2.8^{\circ}$. NeuralGCM employs a dynamical core to solve the primitive equations using a semi-implicit time-integration scheme and a spectral method. Physical processes on the other hand are emulated by learned physics through a neural network.
NeuralGCM has been implemented in JAX (Bradbury et al., 2018) and supports automatic differentiation. This enables the computation of gradients with respect to both initial conditions and internal system parameters, facilitating backpropagation through the physical dynamics and neural network components. In this work, we compute gradients only with respect to the initial variables involved in the dynamical core of NeuralGCM, keeping all other parameters fixed. The loss, as defined in Eq. 3, is minimized using gradient descent, specifically with the Adam optimizer from Optax (DeepMind et al., 2020). The optimal perturbations are applied on the spherical harmonic coefficients representation of the variables. We choose NeuralGCM over other possible models because, it has demonstrated competitive forecast skill for temperatures up to 10-day lead times, contains a dynamical core and relies on a single initial condition.
Although the model runs efficiently on a single GPU with relatively low memory requirements, gradient computation demands substantial memory, scaling rapidly with the number of time steps. To address this, we employ gradient checkpointing and chunking strategies to manage memory usage. These techniques store only essential intermediate values during the forward pass, recomputing them during the backward pass to reduce memory overhead (Kochkov et al., 2024). The optimization scheme on the $2.8^{\circ}$ model runs on a 16GB A4000 Nvidia GPU, whereas the $1.4^{\circ}$ model necessitates a 40GB A100 Nvidia GPU.
We investigate extreme events by perturbing the initial conditions primarily around PN2021 using data from the ERA5 reanalysis dataset (Hersbach et al., 2020). We conducted two simulations; their configurations—including the learning rate ($\alpha$), loss-function weights ($\beta,\vec{\lambda}$), forecast lead times, initialization dates, and number of steps are detailed in Table 1.
$\alpha$ $\beta$ $\lambda_{T}$ $\lambda_{SP}$ $\lambda_{\delta}$ $\lambda_{\zeta}$ $\lambda_{SH}$ $\lambda_{SCIWC}$ $\lambda_{SCLWC}$ Initial Date $\tau$ Total integration time steps $10^{-9}$ 20 200 20 2000 2000 200 20 20 June 21st 2021 5 days 11 days 50 $10^{-9}$ 10 100 10 1000 1000 100 10 10 June 21st 2021 5 days 11 days 75
Table 1: Parameters used during optimization process.
The optimized simulations are compared to an ensemble run of the event using the stochastic version of NeuralGCM. This ensemble consists of 75 members. Unlike our approach, which perturbs the initial conditions (inputs to the model), the stochastic model introduces perturbations within the learned physics module. As a result, the perturbations are effectively introduced one time step apart. Additionally, our method perturbs surface pressure, which is not perturbed in the stochastic model. More details about the stochastic model can be found in Kochkov et al. (2024).
1.3 Heatwave Intensity
@@
We define a heatwave event as a period during which the daily temperature exceeds the 99th percentile threshold for consecutive days (Comeau et al., 2023). The length of the event is given by the number of consecutive days the temperature remains above this threshold. The intensity of the heatwave is measured by the average exceedance of the temperature above the threshold over the duration of the event. Specifically, if $L$ denotes the length of the event, $T_{i}$ the mean temperature timeseries over a region, and $T_{\text{thresh}}$ the 99th percentile threshold, then the intensity $I$ is defined as

### v3

@@ Abstract
The 2021 Pacific Northwest (PN2021) heatwave shattered historical temperature records, culminating in Lytton, Canada’s unprecedented 49.6C observation—a 4.6C increase over the country’s previous record measurement (White2023; ThePacificNorthwestHeatWave). This event, virtually implausible under preindustrial conditions (esd-13-1689-2022), exemplifies a critical challenge in climate science: determining the upper bounds of what is physically possible for different weather extremes under current or future climatic conditions.
The PN2021 heatwave emerged from persistent atmospheric blocking sustained by large-scale Rossby waves that disrupted zonal flow and stalled a high-pressure system over the region (ThePacificNorthwestHeatWave; White2023). This large-scale setup was fueled by upstream dynamics. Mo2022 linked it to anomalous atmospheric river activity, while https://doi.org/10.1029/2021GL097699 identified that diabatic heating within the warm conveyor belt of an upstream cyclone provided the necessary Rossby wave activity to establish the block. Once established, the block suppressed cloud formation and drove prolonged subsidence, adiabatically warming near-surface air masses (MeteorologicalAnalysisofthePacificNorthwestJune2021Heatwave). White2023 corroborated the importance of these mechanisms and estimated via four-day backward trajectory analysis that diabatic processes accounted for approximately 78% of the net temperature change of air parcels entering the region, with the remaining $\sim 22\%$ attributed to adiabatic warming from subsidence. Locally, dry soil conditions further intensified these temperatures through non-linear land-atmosphere interactions (Bartusek2022; TheInfluenceofSoilMoistureontheHistoric2021PacificNorthwestHeatwave; https://doi.org/10.1029/2022EF002967). By studying a 100-member ensemble of PN2021 with varying initial land surface conditions, https://doi.org/10.1029/2025EF006216 found that variations in antecedent soil moisture led to a spread of approximately $3^{\circ}$C in peak temperatures, largely driven by regions shifting into a transitional evaporation regime where latent heat flux becomes highly sensitive to soil moisture.
@@
To systematically explore such extremes, storylines are increasingly used, representing physically consistent sequences of weather events that depict how a counterfactual extreme event might occur (Hazeleger2015; Shepherd2019; https://doi.org/10.1029/2020EF001783). This approach enables a mechanistic exploration of how minor perturbations can lead to the amplification of extreme events. Here, we use a novel differentiable modeling framework to demonstrate that targeted initial-condition perturbations can further amplify these typical extreme trajectories, giving extreme heatwave storylines.
@@
Identifying storylines for the most extreme weather events is a needle-in-a-haystack problem due to their inherent rarity. The traditional approach is the use of single-model initial-condition large ensembles (Deser2020; Suarez-Gutierrez2020; esd-12-401-2021; Diffenbaugh_2021), and more recently, so-called huge ensembles (mahesh2024hugeensemblesidesign; mahesh2024hugeensemblesiiproperties), in which vast numbers of model runs allow the exploration of a wide range of potential outcomes. By systematically increasing ensemble size, the chances of capturing low-probability extremes increase. However, these ensembles are computationally demanding and not very effective at sampling the full range of outcomes. In addition, due to their high computational cost, it is virtually impossible to perform such ensembles using kilometer-scale simulations, which are required to well simulate some types of weather extreme events (e.g., extreme convective precipitation).
@@
In recent years, a number of approaches have been proposed to generate extreme event storylines (doi:10.1073/pnas.1712645115; https://doi.org/10.1029/2018MS001419; 10.1063/1.5081461; gmd-13-763-2020; VeryRareHeatExtremesQuantifying; Fischer2023). These methods focus computing resources on specific extreme events, instead of continuous long simulations. Some approaches enhance the likelihood of simulating extreme events by constructing targeted ensembles (doi:10.1073/pnas.1712645115; 10.1063/1.5081461; Fischer2023). Fischer2023 focus on generating an initial condition ensemble of climate model simulations of known extreme events using a method named ensemble boosting. They applied this approach to the PN2021 heatwave and by perturbing the initial conditions using numerical noise for 500 members, they found a 5-day running average of daily maximum temperature anomalies up to 2.9C larger than the unperturbed event. Other approaches to construct storylines of extreme events use the large deviation algorithm (doi:10.1073/pnas.1712645115; https://doi.org/10.1029/2020GL091197; Statisticalanddynamicalaspectsofextremelyhotsummersin) where an ensemble of simulations is ran and members are periodically pruned or cloned such that an ensemble most likely to lead to an extreme event is generated. Focusing on western European heatwaves and using the large deviation algorithm, doi:10.1073/pnas.1712645115 generated an ensemble which has a mean 2C anomaly compared to a control ensemble of 128 members. Other applications of the algorithm showed its ability to identify even more extreme events, with ensembles with mean anomalies of 4C (https://doi.org/10.1029/2020GL091197). Meanwhile, https://doi.org/10.1029/2018MS001419 introduced a variational data assimilation technique, optimized with a 4D-Var inspired method to intensify past extreme tropical cyclones with minimal perturbations. This approach is closely related to the method presented here; however, we leverage automatic differentiation and computationally efficient ML-based models.
@@
Recent advances in machine learning (ML) have led to the development of transformative tools for weather and climate modeling. Neural network architectures like GraphCast (lam2023graphcastlearningskillfulmediumrange), Pangu-Weather (Bi2023), FourCastNet (pathak2022fourcastnetglobaldatadrivenhighresolution; 10.1145/3592979.3593412), and FuXi (Chen2023) have demonstrated forecasting skill comparable to that of traditional numerical weather prediction systems, but at significantly reduced computational costs (https://doi.org/10.1029/2023MS004019; ValidatingDeepLearningWeatherForecasts; ennis2025turningheatassessing2m; zhang2025numericalmodelsoutperformai). In addition to their reduced computational costs, these models by construction allow us to define optimization problems on them that can be solved through gradient-based optimizers. https://doi.org/10.1029/2023MS004019 introduced a standardized benchmark to compare the various ML models against ERA5 and the European Centre for Medium-Range Weather Forecasts’s (ECMWF) integrated forecast system (IFS). Using this benchmark, it is shown that deterministic, data-driven methods such as Pangu-Weather, GraphCast, and FuXi result in similar root-mean-square error (RMSE) in forecasting near-surface temperature, wind, and pressure up to ten days ahead. However, their forecast skill deteriorates rapidly for longer lead times, resulting in overly smoothed predictions. Using three case studies, ValidatingDeepLearningWeatherForecasts evaluated GraphCast, Pangu-Weather and FourCastNet against ERA5 reanalysis and ECMWF’s IFS for the PN2021 heatwave, the 2023 South Asian humid heatwave and a 2021 North American winter storm. They find that all data-driven models systematically underestimate the peak 2m temperature during the PN2021 heatwave, with root mean square error (RMSE) values at grid points near Vancouver, Seattle, and Portland exceeding twice the 10-day IFS error and reaching up to four times that value in Portland. During the South Asian humid heatwave, data-driven forecasts of heat index computed from 2-m air temperature and 1000-hPa relative humidity underpredicted observed peaks more strongly than IFS, particularly over Bangladesh. For the North American winter storm, data-driven forecasts of wind chill at College Station, Texas, achieved lower peak errors than IFS, with Pangu-Weather and GraphCast outperforming the operational model.
@@
An alternative to purely data-driven approaches is the use of hybrid models, such as NeuralGCM (Kochkov2024), which combines a traditional dynamical core with ML components. (duan2024aibasedclimatemodelevaluation) have shown NeuralGCM’s ability to hindcast the PN2021 heatwave, though due to the lack of processes (such as land-atmosphere feedbacks), the intensity of the heatwaves tends to be underestimated. Similarly to the purely data-driven models, NeuralGCM produces surface variables forecasts with skill comparable to that of the ECMWF IFS system https://doi.org/10.1029/2023MS004019. Moreover, the use of the dynamical core both prevents the evolved fields from being overly smoothed and enhances numerical stability. These benefits allow for longer time integrations and make the model suitable for climate studies (Kochkov2024).
@@
These new types of models are by construction differentiable through automatic differentiation (gmd-16-3123-2023). The automatic differentiation property enables efficient optimization, allowing gradient-based exploration storylines in high-dimensional climate models. This is in line with many new extreme event opportunities enabled by ML models (https://doi.org/10.1002/wcc.914; Camps-Valls2025). Leveraging automatic differentiation, recent studies have implemented variational data assimilation techniques using neural networks, with applications ranging from toy models, such as the Lorenz 96 system (75462), to reduced-order physical representations of the atmosphere (solvik20244dvarusinghessianapproximation; manshausen2024generativedataassimilationsparse). Additionally, vonich2024predictabilitylimit2021pacific demonstrated that the differentiability of GraphCast allows for a more accurate reconstruction of the initial conditions that led to the PN2021 heatwave compared to using ERA5 reanalysis data. Baño-Medina2025 explores the use of ML models and automatic differentiation to perform sensitivity analysis of the initial conditions leading to the development of cyclone Xynthia. Their findings suggest that gradients computed from the data-driven weather model at a 36-hour lead time exhibit sensitivity structures that closely resemble those generated by the adjoint of a dynamical model. In other words, the evolved perturbations from both approaches lead to similar impacts on the cyclone’s evolution.
@@
In this study, we focus on the PN2021 heatwave event due to its well-documented synoptic drivers, and its prevalence in extreme event studies (Lucarini_2023; Fischer2023; White2023; esd-13-1689-2022). We use the automatic differentiation feature of the NeuralGCM model to optimize perturbed initial conditions, and we identify trajectories where enhanced geopotential height anomalies intensify downstream near-surface temperature extremes. These storylines reveal heatwave intensity increases of $3.7^{\circ}$C beyond the extreme temperatures obtained from a 75-member ensemble run using NeuralGCM for the event, analogous to the ensemble boosting approach (Fischer2023). Our results demonstrate the potential of differentiable hybrid models for investigating worst-case scenarios, offering a computationally efficient alternative to traditional, computationally expensive, large ensembles.
@@ 1.1 Initial Conditions Optimization Problem
Our goal is to find the worst-case physically plausible heatwave trajectory our model can produce. To achieve this, we must find the specific, small perturbations to a known initial state that will evolve into the most extreme event. This search is formulated as an optimization problem, where we define a loss function that the model will automatically minimize in an iterative way to find these optimal initial-state perturbations. Formally, a suitable loss function for our problem is one that
  1. 1.
    maximizes a target extreme event, and
  2. 2.
    minimizes the introduced perturbation.
@@
with $b$ representing some nonlinear operator, $t$ the time and $\vec{x}_{0}$ some initial state $t_{0}$. The solution is given by $\vec{X}(t)=S_{t}\vec{x}_{0}$ where $S_{t}$ is the evolution operator up to time $t$. The core aim of the optimization problem is to identify the initial conditions $\vec{x}_{0}$ that drive the dynamical system toward an extreme desired state, represented by a target observable $\mathcal{O}(\vec{X}(t))$. Given a baseline initial state $\vec{x}_{0}^{b}$, we define a perturbation $\Delta\vec{x}_{0}=\vec{x}_{0}^{b}-\vec{x}_{0}$. The optimization problem is formulated in terms of minimizing the following loss function:
@@
In particular, we pick our observable $(\mathcal{O}(\vec{X}(t)))$ to be the temperature over a domain $\mathcal{D}$ and over a period of time $\tau$ at the 1000-hPa pressure level of the model ($\frac{1}{\tau|D|}\int_{0}^{\tau}\int_{D}T_{1000}(\phi,\theta,t)\,d\phi\,d\theta\,dt$), with $\phi,\theta$ being the longitude and latitude. Multiple functions $F(\mathcal{O}(\vec{X}$) can be considered, but our main results uses $F(X)=\frac{c}{X}$ which gives us the loss :
$\mathcal{L}(T_{1000},\Delta x_{0})=\underbrace{\beta\frac{T_{\text{ref}}}{\frac{1}{\tau|D|}\int_{0}^{\tau}\int_{D}T_{1000}(\phi,\theta,t)\,d\phi\,d\theta\,dt}}_{\text{Temperature objective term}}+\underbrace{\sum_{i}\lambda_{i}\frac{(\Delta x_{0,i})^{2}}{(\Delta x_{\text{ref},i})^{2}}}_{\text{Perturbation penalty term}}$ (3)
@@
where $\mathcal{D}$ corresponds to the region shown in Fig. 4, and $\tau$ is set to 5 days. This 5-day period was chosen to fully encompass the three peak days of the PN2021 event, with a two-day buffer at the end, which we found aided in optimization. The terms in the loss function are normalized by their initial means, with $T_{\text{ref}}$ representing a characteristic temperature scale and $\Delta x_{\text{ref},i}$ denoting a reference perturbation scale for each perturbed variable $i=\{$Temperature, Surface Pressure, Vorticity, Divergence, Specific Humidity, Specific Cloud Ice Water Content, Specific Cloud Liquid Water Content}\}. The normalization scale for each perturbed variable, $\Delta x_{\text{ref},i}$, is defined as the absolute mean of each respective initial field. This ensures each terms is of a similar magnitudes.
@@
Once the simulation are optimized, we evaluate their success through an intensity metric for the heatwaves. We define a heatwave event as a period during which the daily temperature exceeds the 99th percentile threshold for consecutive days (SP3/Z4Y0LK_2023). This definition relies on the persistence of temperature extremes (see also heatwave intensity definition); if the temperature drops below the threshold for even a single day, the event is considered terminated, and any subsequent exceedances are treated as distinct, separate events. The intensity of the heatwave is measured by the average exceedance of the temperature above the threshold over the duration of the event. Specifically, if $L$ denotes the length of the event, $T_{i}$ the mean temperature timeseries over a region, and $T_{\text{thresh}}$ the 99th percentile threshold, then the intensity $I$ is defined as
@@
1.2 Numerical Implementation using NeuralGCM
To simulate the dynamics and evaluate the loss function, we use the NeuralGCM model (Kochkov2024). Most of the experiments are performed with a horizontal grid spacing of $2.8^{\circ}$ (denoted as NeuralGCM2.8) because it is more computationally tractable and a climate simulation is available at this resolution. For sensitivity analysis we also consider simulations performed using a horizontal grid spacing of $1.4^{\circ}$. NeuralGCM employs a dynamical core to solve the primitive equations using a semi-implicit time-integration scheme and a spectral method. Physical processes on the other hand are emulated by learned physics through a neural network.
NeuralGCM has been implemented in JAX (jax2018github) and supports automatic differentiation. This enables the computation of gradients with respect to both initial conditions and internal system parameters, facilitating backpropagation through the physical dynamics and neural network components. In this work, we compute gradients only with respect to the initial variables involved in the dynamical core of NeuralGCM, keeping all other parameters fixed. The loss, as defined in Eq. 3, is minimized using gradient descent, specifically with the Adam optimizer from Optax (deepmind2020jax). The optimal perturbations are applied on the spherical harmonic coefficients representation of the variables. We choose NeuralGCM over other possible models because, it has demonstrated competitive forecast skill for temperatures up to 10-day lead times, contains a dynamical core and relies on a single initial condition.
Although the model runs efficiently on a single GPU with relatively low memory requirements, gradient computation demands substantial memory, scaling rapidly with the number of time steps. To address this, we employ gradient checkpointing and chunking strategies to manage memory usage. These techniques store only essential intermediate values during the forward pass, recomputing them during the backward pass to reduce memory overhead (Kochkov2024). The optimization scheme on the $2.8^{\circ}$ model runs on a 16GB A4000 Nvidia GPU, whereas the $1.4^{\circ}$ model necessitates a 40GB A100 Nvidia GPU.
We investigate extreme events by perturbing the initial conditions primarily around PN2021 using data from the ERA5 reanalysis dataset (ERA5). We conducted two independent optimization experiments, hereafter referred to as “EXP50” and “EXP75”. Their configurations—including the learning rate ($\alpha$), loss-function weights ($\beta$, $\lambda_{i}$), forecast lead times, initialization dates, and number of gradient descent steps (N)—are detailed in Table 1. These parameters were selected via an experimental approach analogous to machine learning hyperparameter tuning, as an exhaustive automated search would be computationally prohibitive. We initially selected $N=75$ to establish a baseline comparable in computational cost to a 75-member ensemble. Subsequently, we performed the $N=50$ experiment to assess whether similar results could be achieved with fewer resources. This required retuning the $\lambda_{i}$ parameters; generally, a larger $N$ implies a longer search time, allowing perturbations to grow larger, which in turn necessitates a higher $\lambda$ to constrain their size. Finally, forecast lead times were chosen to strike a balance: sufficiently close to the event to ensure forecastability, yet distant enough to allow the introduced perturbations adequate time to evolve.
Experiment name $\alpha$ $\beta$ $\lambda_{T}$ $\lambda_{SP}$ $\lambda_{\delta}$ $\lambda_{\zeta}$ $\lambda_{SH}$ $\lambda_{SCIWC}$ $\lambda_{SCLWC}$ Initial Date $\tau$ Total integration time steps EXP50 $10^{-9}$ 20 200 20 2000 2000 200 20 20 June 21st 2021 5 days 11 days 50 EXP75 $^{-9} 10 10$0 10 1000 1000 100 10 10 June 21st 2021 5 days 11 days 75
Table 1: Parameters used during the optimization process. Each row corresponds to one experiment. The coefficients $\lambda_{T}$, $\lambda_{SP}$, $\lambda_{\delta}$, $\lambda_{\zeta}$, $\lambda_{SH}$, $\lambda_{SCIWC}$, and $\lambda_{SCLWC}$ control the relative weight of the temperature term, the surface pressure term, the divergence term, the vorticity term, the specific humidity term, and the ice and liquid cloud water terms in the loss function. The parameter $\beta$ sets the strength of the temperature objective term. The number of iteration steps differs between the two experiments in order to explore the effect of longer and shorter optimization procedures while all other settings are kept fixed. The quantity $\tau$ denotes the forecast lead time used when computing the loss.
The optimized simulations are compared to an ensemble run of the event using the stochastic version of NeuralGCM. This ensemble consists of 75 members. Unlike our approach, which perturbs the initial conditions (inputs to the model), the stochastic model introduces perturbations within the learned physics module. As a result, the perturbations are effectively introduced one time step apart. Additionally, our method perturbs surface pressure, which is not perturbed in the stochastic model. More details about the stochastic model can be found in Kochkov2024.